OpenStudy (anonymous):

the limit of x^3lnx as x approaches 0 is indeterminate.evaluate it by using L'Hopital rule

OpenStudy (anonymous):

so wats the question????????? not to sure what u're trying to find... it's x cubed what use the Equation thing

OpenStudy (anonymous):

\[\lim_{x\to0^+}x^3\ln x=0\cdot(-\infty)\\ \lim_{x\to0^+}\frac{\ln x}{\frac{1}{x^3}}=\frac{-\infty}{\infty}\] Now you can use L'Hopital's rule.

OpenStudy (anonymous):

thank you. but then when i use the L'Hopital's rule i get non-stoping \[-\infty/\]

OpenStudy (anonymous):

Applying L'Hopital's rule, you get \[\begin{align*}\lim_{x\to0}\frac{\dfrac{1}{x}}{-\dfrac{3}{x^4}}&=\lim_{x\to0}\dfrac{1}{x}\left(-\dfrac{x^4}{3}\right)\\ &=-\frac{1}{3} \lim_{x\to0}x^3\\ \end{align*}\]

OpenStudy (anonymous):

thanks , i get it now